Carla plans to invest in a property that after 3 years will yield $1200 at the beginning of each month indefinitely. How much should Carla be willing to pay if alternative investment yields 9% compounded monthly?
1
Expert's answer
2020-07-26T17:20:36-0400

Compound interest : A = P (1+ r/100)n

one year has 12 months therefore three years will be 3*12 = 36 months.

amount = 1200

p= ?

n = 36

r = 9 % since compounding is in months r will be 9%/12

1200 = p (1 + 9%/12)36

1200 = p (1.0075)36

1200 = p ( 1.3086)

p = 1200/1.3086

P = 917.01

therefore she will be paying approximately $917.